Solve the below differential equation
y'''+3y''+2y'-5y=sin2t
1
Expert's answer
2019-05-31T10:19:08-0400

Consider the equation


"y'''+3y''+2y'-5y=0"

We write the characteristic equation


"k^3+3\times k^2+2\times k -5=0"

Using the Cardano formulas we get the roots of the equation


"k_1=0.9"

"k_2=-1.95-1.31\times i"

"k_3= -1.95 +1.31\times i"

where


"i=\sqrt {-1}."

Therefore solving the equation


"y'''+3y''+2y'-5y=0"

"y=C_1\times e^ {0.9t} + C_2\times e^ {-1.95t}sin (1.31t)+C_3\times e^ {-1.95t}cos (1.31t)"

where


"C_1, C_2, C_3"

are constants. Assume a partial solution of the equation is


"y_p=Asin(2t)+Bcos(2t)"

where A, B are constants. Hence


"y'_p=2Acos(2t)-2Bsin(2t)"


"y''_p=-4Asin(2t)-4Bcos(2t)"

"y'''_p=-8Acos(2t)+8Bsin(2t)"

Substituting in the equation we get


"(-3A+6B)sin(2t)+(-6A-3B)cos(2t)=sin(2t)."

As result one gets


"A=- \frac {1} {15}"

"B= \frac{2} {15}"

Therefore the solution of the equation


"y'''+3y''+2y'-5y=sin(2t)"

is

"y=C_1\times e^ {0.9t} + C_2\times e^ {-1.95t}sin (1.31t)+C_3\times e^ {-1.95t}cos (1.31t)+"

"- \frac {1} {15} sin(2t) + \frac {2} {15} cos(t)"

where


"C_1, C_2, C_3"

are constants.

Answer.

"y=C_1\times e^ {0.9t} + C_2\times e^ {-1.95t}sin (1.31t)+C_3\times e^ {-1.95t}cos (1.31t)- \frac {1} {15} sin(2t) + \frac {2} {15} cos(t)"

where

"C_1, C_2, C_3"

are constant.